OpenStudy (anonymous):

What is the solution to the equation? z – (–1.5) = –2.3 z =

OpenStudy (anonymous):

z – (–1.5) = –2.3 z+1.5 = -2.3 z= -2.3-1.5 = -3.8 z= -3.8

OpenStudy (anonymous):

since - and - becomes plus

OpenStudy (anonymous):

yes he is right